OpenStudy (anonymous):

Question 2. 2. Which of the following methods would be the easiest to use to solve x2 – 11 = 0? (Points : 1) factoring all three methods would be easy and effective isolating the x2 term and finding the square root of both sides using the quadratic formula

OpenStudy (solomonzelman):

your opinion?

OpenStudy (anonymous):

Isolating the x terms?

OpenStudy (solomonzelman):

yes

OpenStudy (anonymous):

ok thank you I have a few more would you mind helping me with?

OpenStudy (wolf1728):

When I was in algebra class that's what we called the difference of two squares

OpenStudy (solomonzelman):

\(x^2-11=0\) \(x^2-11\color{red}{+11}=0\color{red}{+11}\) \(x^2=11\) \(x=\pm \sqrt{11}\)

OpenStudy (anonymous):

Thank you!

OpenStudy (solomonzelman):

well, I can do the quadratic formula with b=0, no trouble, but would i?

OpenStudy (solomonzelman):

yw.....

OpenStudy (anonymous):

so could you help me with a few more?

OpenStudy (wolf1728):

I'll try one more

OpenStudy (anonymous):

3. Which of the following methods would be the easiest to use to solve x2 – 7 = 0? (Points : 1) all three methods would be easy and effective factoring isolating the x2 term and finding the square root of both sides using the quadratic formula

OpenStudy (anonymous):

Thank you so much you have no idea!

OpenStudy (wolf1728):

Same answer as the previous question.

OpenStudy (anonymous):

ok so none of the others are easier?

OpenStudy (anonymous):

Is this fair to ask because all of them are fine to use?

OpenStudy (wolf1728):

No - and the only difference between the 2 equations is the minus 11 changes to minus 7

OpenStudy (anonymous):

ok thank you!

OpenStudy (wolf1728):

All will give you the answer but I think the easiest is " isolating the x2 term , etc

OpenStudy (anonymous):

ok great thank you! I will try them because I really dont think it is fair to ask because all of them are pretty easy to use as well it is all a personal chose.

OpenStudy (wolf1728):

Okay - glad to help :-)